IDENTIFICATION OF THE ^{3}Σ_{u}^{−} AND ^{3}Π_{u} RYDBERG STATES OF O_{2} BY DETECTING O(^{1}D)

\underline{J. B. NEE} AND P. C. LEE, DEPARTMENT OF PHYSICS AND CHEMISTRY, NATIONAL CEN-TRAL UNIVERSITY, CHUNG-LI, TAIWAN, 32054, ROC.

The excited states of O_{2} in 100-135 nm region consists of many Rydberg states which are still not well understood. We have investigated the excited states of O_{2} dissociated into O(^{1}D)+O(^{3}P). The experiment was conducted in a gas celll using the synchrotron radiation as the excitation source. O(^{1}D) was detected by the infrared emission at 762 nm produced from the O_{2}(b^{1}Σ_{g}^{+}) formed by the energy transfer process from O(^{1}D) to O_{2}. A spectrum between 105 and 130 nm for producing O(^{1}D) was obtained, which revealed the vibrational progression of the n^{3}Σ_{u}^{−} states for n=2 to 4. By comparing with the absorption spectrum, the ^{3}Π_{u} state correlated with O(^{3}P)+O(^{3}P) was also obtained.